### v1

# Some remarks on L-equivalence for cubic fourfolds and hyper-Kähler manifolds

Abstract
We prove that if two very general cubic fourfolds are L-equivalent then they are isomorphic, and we observe that there exist special cubic fourfolds which are L-equivalent but not isomorphic. As a consequence, we provide examples of hyper-Kähler manifolds which are L-equivalent but not birational. We also provide further examples in support of the fact that L-equivalent hyper-Kähler manifolds should be D-equivalent, as conjectured by Meinsma.
1 Introduction
1.1 L-equivalence
The Grothendieck ring of varieties $K_{0}(\operatorname{Var}_{k})$ over a field $k$ is the free abelian group generated by isomorphism classes of varieties over $k$, modulo the relations
$X]=[X\setminus Z]-[Z]$
for every closed subvariety $Z\subset X$. The ring structure is given by the product
$X]\cdot[Y]:=[X\times_{k}Y],$
which is commutative with unit the class of a point $\operatorname{Spec}(k)]$. The class of the affine line
$\mathbb{L}:=[\mathbb{A}^{1}_{k}]\in K_{0}(\operatorname{Var}_{k})$
is called the Lefschetz motive. The class of a variety in the Grothendieck ring of varieties is a motivic invariant, it also carries information about the birational geometry of the variety and about its derived category.
It was proved by Larsen and Lunts [LarsenLunts:motivic] that the quotient $K_{0}(\operatorname{Var}_{k})/\mathbb{L}$ is isomorphic to the free abelian group generated by the stably birational classes of smooth projective varieties. Two varieties $X,Y$ are called D-equivalent if there is an equivalence $D^{b}(X)\cong D^{b}(Y)$ of their bounded derived categories of coherent sheaves. If two varieties are D-equivalent then results of Kawamata [kawamata2002d] and Kontsevich [Kontsevich:motivic_integration] prove that their classes are equal in the completion (with respect to the dimension filtration) of the localization $K_{0}(\operatorname{Var}_{k})[\mathbb{L}^{-1}]$. It is then interesting to study the localization $K_{0}(\operatorname{Var}_{k})\to K_{0}(\operatorname{Var}_{k})[\mathbb{L}^{-1}]$, with the difficulty that Borisov proved [Borisov:class_of_affine_line] that the class $\mathbb{L}$ is a zero-divisor in $K_{0}(\operatorname{Var}_{\mathbb{C}})$. This lead Kuznetsov and Shinder to the following definition.
Definition 1.1 ([kuznetsov2018grothendieck]).
Two varieties $X$ and $Y$ over $k$ are L-equivalent if there exists a positive integer $r$ such that
$\mathbb{L}^{r}\cdot([X]-[Y])=0\in K_{0}(\operatorname{Var}_{k}).$
It was conjectured by Kuznetsov–Shinder [kuznetsov2018grothendieck, Conjecture 1.6] that simply connected algebraic varieties that are D-equivalent should be also L-equivalent. By a result of Orlov, if two varieties with ample canonical class or anticanonical class are D-equivalent then they are isomorphic. In particular, they are clearly L-equivalent. On the other hand, when the canonical bundle is trivial, this implication is proved to be false in [meinsma2025counterexamples, Theorem 1.3]. Indeed, there exist examples of projective hyper-Käher manifolds that are D-equivalent but not L-equivalent. The opposite implication is not entirely clear in general, and Meinsma conjectures [meinsma2025counterexamples, Conjecture 1.5] that two L-equivalent projective hyper-Kähler manifolds must be D-equivalent.
1.2 Results
We will always assume $k=\mathbb{C}$. The results we present are mainly based on the Hodge-theoretical and lattice-theoretical implications of L-equivalence, which were developed by Efimov [efimov2018some] and Meinsma [meinsma2024equivalence].
Our first contribution is to determine wether L-equivalent cubic fourfolds need to be isomorphic (which is the same as being D-equivalent).
Theorem 1.2 (Theorem 3.3).
Let $Y\subset\mathbb{P}^{5}$ be a very general cubic fourfold. If $Y^{\prime}$ is a cubic fourfold which is L-equivalent to $Y$, then $Y\cong Y^{\prime}$.
We say that a cubic fourfold $Y$ is special if it contains a surface $S$ which is not homologous to the square of a hyperplane class $h^{2}=\eta_{Y}\in A(Y):=H^{4}(Y,\mathbb{Z})\cap H^{2,2}(Y)$, i.e. the algebraic lattice $A(Y)$ has rank at least $2$. Special cubic fourfolds belong to Hassett divisors $\mathcal{C}_{d}\subset\mathcal{C}$ in the moduli space of cubic fourfolds with $d\equiv 0,2\mod 6$ and $d>6$, see [hassett2000special]. A very general cubic fourfold in the moduli space of cubic fourfolds $\mathcal{C}$ is not special, we will sometimes refer to cubic fourfolds that are very general in a Hassett divisor $\mathcal{C}_{d}$ so that they do not belong to any other divisor $\mathcal{C}_{d^{\prime}}$ with $d^{\prime}\not=d$.
We observe that Theorem 1.2 might be not true when considering a cubic fourfold which is not very general.
Remark 1.3 (3.5).
There exist special cubic fourfolds $Y$ and $Y^{\prime}$ which are L-equivalent, but not isomorphic (and in particular not D-equivalent). [^3] Similar general results about L-equivalent cubic fourfolds are independently achieved by Meinsma and Moschetti in a paper to appear, where further examples of this phenomenon are studied.
Such examples are cubic fourfolds $Y,Y^{\prime}\in\mathcal{C}_{20}$ which were studied by Fan-Lai [Fan_Lai:New_rational], they are birational to each other. These cubic fourfolds are also Furier-Mukai (FM) partners, meaning that there is a Furier-Mukai equivalence $\mathcal{A}_{Y}\cong\mathcal{A}_{Y^{\prime}}$ between the Kuznetsov components [kuznetsov2009derived], here $\mathcal{A}_{Y}:=\langle\mathcal{O},\mathcal{O}(1),\mathcal{O}(2)\rangle^{\perp}\subset D^{b}(Y)$.
Theorem 1.4 (Theorem 3.6).
Consider a very general cubic fourfold $Y\in\mathcal{C}_{d}$ with $d\equiv 0,2\mod 6$ and $d$ not divisible by $9$. If $Y^{\prime}$ is a cubic fourfold which is L-equivalent to $Y$, then $Y$ and $Y^{\prime}$ are FM partners.
We then wonder if this is always true and pose the following.
Conjecture 1.5.
If $Y$ and $Y^{\prime}$ are L-equivalent cubic fourfolds then they are FM partners.
We now turn to the case of hyper-Kähler manifolds. The first examples of hyper-Kähler manifolds that are L-equivalent and D-equivalent but not birational were constructed by Okawa in [okawa2021example]. Here we give further examples of L-equivalent hyper-Kähler manifolds which are not birational, we expect them to be also D-equivalent in view of [meinsma2025counterexamples, Conjecture 1.5]. Recall that the Fano variety of lines $F(Y)$ of a cubic fourfold $Y$ is hyper-Kähler deformation equivalent to the Hilbert square of a K3 surface [beauville1985variete].
Theorem 1.6 (3.5 and 4.5).
Let $Y\in\mathcal{C}_{20}$ be a general cubic fourfold, then there exists a cubic fourfolds $Y^{\prime}$ such that the Fano varieties of lines $F(Y)$ and $F(Y^{\prime})$ are L-equivalent but not birational.
We also provide some additional evidence for the Meinsma’s conjecture, in some particular cases. If $X$ is hyper-Kähler, we consider the transcendental lattice $T(X)\subset H^{2}(X,\mathbb{Z})$, the smallest sublattice such that $H^{2,0}(X)\subset T(X)\otimes\mathbb{C}$, and the Néron-Severi lattice $\operatorname{NS}(X)\subset H^{2}(X,\mathbb{Z})$ which can be seen as the orthogonal to $\operatorname{NS}(X)$ with respect to the Beauville–Bogomolov–Fujiki form, and it consists of algebraic classes.
Theorem 1.7 (4.6, 4.11, 4.13, 4.14, 4.16).
Let $X$ and $X^{\prime}$ be projective hyper-Kähler manifolds which are deformation equivalent and L-equivalent. Assume that $\operatorname{End}(T(X))=\mathbb{Z}$, then:
  1. 1.
    If $X$ and $X^{\prime}$ are birational to the (twisted) LSV manifolds associated to very general cubic fourfolds, then $X$ and $X^{\prime}$ are birational.
  2. 2.
    If $X$ and $X^{\prime}$ are moduli spaces on the Kuznetsov component of very general cubic fourfolds $Y$ and $Y^{\prime}$, then $Y\simeq Y^{\prime}$ and $X$ is birational to $X^{\prime}$.
  3. 3.
    If $X$ is of OG10 type, $\operatorname{NS}(X)=U\oplus A_{2}(-1)$ and $\operatorname{disc}T(X)=1$, then $X$ and $X^{\prime}$ are birational and D-equivalent.
  4. 4.
    If $X=S^{[n]}$ for n1n-1 a prime power and a K3 surface $S$ with $U\subseteq\operatorname{NS}(S)$ and $\operatorname{rk}\operatorname{NS}(S)\not=4$, then in case $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})=m$ is coprime to 2(n1)2(n-1) we have that $X$ and $X^{\prime}$ are D-equivalent.
Here $\operatorname{disc}T(X)$ denotes the discriminant of $T(X)$ as a lattice and $\operatorname{End}(T(X))$ denotes the algebra of Hodge endomorphisms. We remark that the hypothesis $\operatorname{End}(T(X))=\mathbb{Z}$ is satisfied by a general element of a lattice-polarized family, see 2.8.
Outline
We collect in § 2 preliminaries and known results about lattices, Hodge structures, and L-equivalence. In § 3 we present our results about cubic fourfolds. The subsections of § 4 are dedicated to various constructions of hyper-Kähler manifolds involving cubic fourfolds and K3 surfaces.
Acknowledgements
The discussions about cubic fourfolds and the collaboration started at the IX Japanese-European Symposium on Symplectic Varieties and Moduli Spaces. We would like to thank the organizers for the nice environment of collaboration offered during the conference.
Simone Billi would like to thank the Matthias Schütt and the Riemann Center for Geometry and Physics of Hannover for hosting and supporting him during the writing of this article.
2 Hodge lattices and L-equivalence
2.1 Hodge lattices
A lattice is a finitely generated free abelian group $L$ with a non-degenerate symmetric bilinear form $b\colon L\times L\to\mathbb{Z}$. For elements $v,w\in L$ we will write $v^{2}:=b(v,v)$ and $v\cdot w:=b(v,w)$. A lattice $L$ is called even if for any $v\in L$ we have $v^{2}\in 2\mathbb{Z}$, it is called odd otherwise. We denote by $O(L)$ the group of isometries, i.e. isomorphisms preserving the bilinear form. We consider the dual lattice $L^{\vee}:=\operatorname{Hom}(L,\mathbb{Z})\cong L\otimes_{\mathbb{Z}}\mathbb{Q}$, which naturally contains $L$, and let
$D(L):=L^{\vee}/L$
be the discriminant group. It is a torsion group endowed with a symmetric bilinear form, which takes values in $\mathbb{Q}/2\mathbb{Z}$ when $L$ is even. We moreover let
$\operatorname{disc}(L):=|D(L)|\in\mathbb{Z}$
be the discriminant of $L$, which is also equal to the absolute value of the determinant of a Gram matrix for $b$. The length of the discriminant group $l(D(L))$ is the smaller number of generators for $D(L)$, it does not exceed the rank of $L$. A lattice $L$ is called unimodular if $\operatorname{disc}(L)=1$, or equivalently $L=L^{\vee}$. The genus of a lattice is the data of its signature, parity and discriminant form. The genus is clearly an isometry invariant, but there are non-isometric lattices with the same genus.
For an integer $n\in\mathbb{Z}$ we denote by $L(n)$ the lattice obtained by $L$ multiplying the bilinear form by $n$. We can allow also rational numbers $q\in\mathbb{Q}$ and consider $L(q)$, whose bilinear form takes rational values a priori, and it is a lattice if its quadratic form takes integer values.
An injective morphism of lattices $L\hookrightarrow M$ is called an embedding, in this case we can see $L\subset M$ as a sublattice of $M$ and denote by $L^{\perp}\subset M$ its orthogonal complement. An embedding of lattices $L\subset M$ is called primitive if the quotient $M/L$ is torsion-free, it is called of finite index if the quotient $M/L$ is finite and in this case $M$ is called an overlattice of $L$.
Following [Nikulin_integral], a primitive embedding $L\subset M$ determines a finite index embedding $L\oplus L^{\perp}\subseteq M$ and the group
$G_{L,M}=\frac{M}{L\oplus L^{\perp}}$
is embedded isometrically in $D(L)$ and anti-isometrically in $D({L^{\perp}})$. The group $G_{L,M}$ is called the gluing subgroup. It follows that
$|G_{L,M}|^{2}=\frac{\operatorname{disc}(L)\cdot\operatorname{disc}(L^{\perp})}{\operatorname{disc}(M)}.$ (1)
Lemma 2.1.
Let $L\subseteq\Lambda$ be a primitive sublattice of a unimodular lattice $\Lambda$. Let $T$ be a primitive sublattice of $L$, assume there exist $n\in\mathbb{Z}\setminus\{0\}$ such that $T(n)$ admits a primitive embedding in $L$. If $2\operatorname{rk}T>\operatorname{rk}\Lambda$ then $n=\pm 1$.
Proof.
Assume by contradiction that $n\not=\pm 1$. Then D(T(n))D({T(n)}) contains a subgroup isomorphic to $(\mathbb{Z}/n\mathbb{Z})^{\operatorname{rk}T}$, generated by an integral basis of $T(n)$ multiplied by $\frac{1}{n}$. Since $\Lambda$ is unimodular, from (1) we have that the primitive embedding $T(n)\subset\Lambda$ induces an isomorphism $D({T(n)})\cong D(N)$, where $N=T(n)^{\perp}\subset\Lambda$. This means that $\operatorname{rk}T\leq\operatorname{rk}N=\operatorname{rk}\Lambda-\operatorname{rk}T$ which implies $2\operatorname{rk}T\leq\operatorname{rk}\Lambda$, a contradiction. ∎
Definition 2.2.
  1. 1)
    A polarized torsion-free integral Hodge structure is called a Hodge lattice.
  2. 2)
    A Hodge structure $H$ of weight $2n$ is called of K3 type if it is torsion free, $\dim_{\mathbb{C}}H^{n+1,n-1}=1$ and $H^{n+i,n-i}=0$ for $i=2,\dots,n$.
  3. 3)
    A Hodge lattice of K3 type is called irreducible if it admits no primitive proper Hodge sublattice of K3 type.
The polarization of a torsion-free Hodge structure gives a lattice structure which is compatible with the Hodge structure. In particular, a polarized Hodge structure of K3 type has the structure of a Hodge lattice. The main example of irreducible Hodge lattice will be the transcendental part of a Hodge structure of K3 type. For a Hodge lattice $H$, we denote by $\operatorname{End}(H)$ the algebra of Hodge endomorfisms, by $\operatorname{Aut}(H)$ the group of Hodge automorfisms and by $O_{Hdg}(H)$ the group of Hodge isometries.
The following very useful and technical result about Hodge lattices will be fundamental.
Proposition 2.3 ([meinsma2024equivalence, Proposition 1.1 and Lemma 3.1]).
Let $T$ and $T^{\prime}$ be irreducible Hodge lattices of K3 type. Assume that $\operatorname{End}(T)=\mathbb{Z}$ and that there is an isomorphism of Hodge structures $T\cong T^{\prime}$. Then there exists $q\in\mathbb{Q}$ such that $T^{\prime}(q)$ is integral and there is a Hodge isometry $T\cong T^{\prime}(q)$.
If $H$ is a polarizable Hodge structure of K3 type and weight $2n$, we consider the algebraic part $A(H):=H^{n,n}\cap H$. Let $T(H)$ be the minimal integral Hodge sub-structure of $H$ such that
$\bigoplus_{i=1}^{n}H^{n+i,n-i}\subset T(H)_{\mathbb{C}},$
we call it the transcendental part. We also set
$G(H):=\frac{H}{A(H)\oplus T(H)}.$
After fixing a polarization, we can consider the orthogonal complement $A(H)^{\perp}$, in this case $(A(H)^{\perp})^{n+i,n-i}=H^{n+i,n-i}$ for any $i=1,\dots,n$ so that $A(H)^{\perp}=T(H)$ if $H$ is torsion-free. If $H$ is torsion free and polarized, then it has the structure of a Hodge lattice. In particular, $A(H)$ and $T(H)$ are primitive sublattices and we get a finite index Hodge embedding $A(H)\oplus T(H)\subseteq H$, so that $G(H)=G_{A(H),H}$ is the gluing subgroup.
We will be mainly interested in the case of Hodge structures of K3 type.
Lemma 2.4.
Let $H$ and $H^{\prime}$ be polarized Hodge structures of K3 type of weight $2n$, assume that $H$ and $H^{\prime}$ are isometric as lattices and that there is a Hodge isometry $f\colon T(H)\cong T(H^{\prime})$. Then there exists a Hodge isometry $h\colon H\cong H^{\prime}$ restricting to $f$ if and only if there is an isometry $g\colon A(H)\cong A(H^{\prime})$ such that $f$ and $g$ induce the same isometry from $G(H)$ to $G(H^{\prime})$.
Proof.
On the lattice-theoretical level, this is precisely [Nikulin_integral, Corollary 1.5.2]. We then observe that such an $h$ is Hodge if and only $f$ is. In fact, since $f=h_{\mid T(H)}$ then $f$ is Hodge from $T(H)$ to $T(H^{\prime})$ if and only if $h$ is, moreover we have $g(A(H)_{\mathbb{C}})=A(H^{\prime})_{\mathbb{C}}$ and also $H_{\mathbb{C}}=A(H)_{\mathbb{C}}\oplus T(H)_{\mathbb{C}}$ since $H$ is polarized (of K3 type), similarly for $H^{\prime}$. ∎
Lemma 2.5.
Let $H$ and $H^{\prime}$ be polarized Hodge structures of K3 type of weight $2n$ such that $H$ and $H^{\prime}$ are isometric as lattices and that $\operatorname{rk}A(H)\geq l(D({A(H)}))+2$. Assume that there is a Hodge isometry $f\colon T(H)\cong T(H^{\prime})$ and that $A(H)$ and $A(H^{\prime})$ are even indefinite and in the same genus. Then there is an isometry $g\colon A(H)\cong A(H^{\prime})$ and a Hodge isometry $h\colon H\cong H^{\prime}$ restricting to $f$ and $g$.
Proof.
By 2.4 it suffices to find an isometry $g\colon A(H)\cong A(H^{\prime})$ inducing the same isometry $G(H)\cong G(H^{\prime})$ as $f$. From the fact that $A(H)$ and $A(H^{\prime})$ are even indefinite and in the same genus with $\operatorname{rk}A(H)\geq l(D({A(H)}))+2$, we know by [Nikulin_integral, Theorem 1.14.2] that they are isometric and that the map $O(A(H))\to O(D({A(H)}))$ is surjective. Hence, we can always find the wanted isometry $g$. ∎
2.2 Implications of L-equivalence
As in [efimov2018some], we consider the additive category $\operatorname{HS}_{\mathbb{Z}}$ of integral polarizable Hodge structures and for any $n$ the subcategory $\operatorname{HS}_{\mathbb{Z},n}$ of integral polarizable Hodge structures of weight $n$. We consider the abelian group $K_{0}(\operatorname{Hdg}_{\mathbb{Z}})$ generated by elements of $\operatorname{Hdg}_{\mathbb{Z}}$ and subject to the relations
$H_{1}\oplus H_{2}]=[H_{1}]+[H_{2}]$
for any $H_{1},H_{2}\in\operatorname{Hdg}_{\mathbb{Z}}$. There is a homomorphism of additive groups
$\operatorname{Hdg}_{\mathbb{Z}}\colon K_{0}(\operatorname{Var}_{\mathbb{C}})\to K_{0}(\operatorname{HS}_{\mathbb{Z}})$
where $\operatorname{Hdg}_{\mathbb{Z}}([X])=[H^{*}(X,\mathbb{Z})]$, that factors through the localization $K_{0}(\operatorname{Var}_{\mathbb{C}})[\mathbb{L}^{-1}]$, see [efimov2018some]. A similar definition can be given for $\operatorname{Hdg}_{\mathbb{Z},n}$.
For the following, we readapt the definitions and results given in [meinsma2024equivalence, Section 4]. Consider an integral polarizable Hodge structure $H\in\operatorname{HS}_{\mathbb{Z},2n}$ of weight $2n$. There is an additive endofunctor $\mathcal{T}\colon\operatorname{Hdg}_{\mathbb{Z},2n}\to\operatorname{Hdg}_{\mathbb{Z},2n}$ associating to a Hodge structure its transcendental part. In particular, if $H]=[H^{\prime}]$ then $T(H)]=[T(H^{\prime})]$. Moreover, we have a group homomorphism
$\mathcal{G}\colon K_{0}(\operatorname{HS}_{\mathbb{Z},2n})\to\mathbb{Q}^{*}$
such that $\mathcal{G}([H])=|G(H)|$.
Recall that we have
$|G(H)|^{2}=\frac{\operatorname{disc}(A(H))\cdot\operatorname{disc}(T(H))}{\operatorname{disc}(H)}.$
Lemma 2.6.
Let $H$ and $H^{\prime}$ be unimodular Hodge lattices of K3 type, and let $T$ and $T^{\prime}$ be their transcendental parts. Assume $H]=[H^{\prime}]\in\operatorname{HS}_{\mathbb{Z}}$, then $\operatorname{disc}(T)=\operatorname{disc}(T^{\prime})$.
Proof.
Since the Hodge lattice $H$ is unimodular, we have $G(H)\cong D(T)$ and hence $\mathcal{G}([H])=|D(T)|=\operatorname{disc}T$. Now, $\operatorname{disc}(T)=\mathcal{G}([H])=\mathcal{G}([H^{\prime}])=\operatorname{disc}(T^{\prime})$. ∎
Lemma 2.7.
Let $T$ and $T^{\prime}$ be Hodge lattices of K3 type. Assume that $\operatorname{End}(T)=\mathbb{Z}$ and that $T]=[T^{\prime}]\in K_{0}(\operatorname{HS}_{\mathbb{Z}})$. If $\operatorname{disc}T=\operatorname{disc}T^{\prime}$ then there is a Hodge isometry $T\cong T^{\prime}(\pm 1)$.
Proof.
Since $\operatorname{End}(T)=\mathbb{Z}$, by [efimov2018some, Theorem 2.3] we have that there exists an isomorphism of Hodge structures $T\cong T^{\prime}$. In this case there is a Hodge isometry $T\cong T^{\prime}(q)$ by 2.3, implying
$\operatorname{disc}T=q^{\operatorname{rk}T}\operatorname{disc}T^{\prime}$
from which we deduce that $q=\pm 1$ since $\operatorname{disc}T=\operatorname{disc}T$. ∎
Remark 2.8.
As a consequence of [geemen2016conjecture, Lemma 9], the condition $\operatorname{End}(T)=\mathbb{Z}$ is satisfied if $T$ is the transcendental part of $H^{4}(Y,\mathbb{Z})$, or $H^{2}(X,\mathbb{Z})$, where $Y$ is a cubic fourfold, and $X$ is a hyper-Kähler manifold, very general in the family polarized by a certain lattice. In particular, this holds if $Y$ is a very general cubic fourfold or $Y$ is very general in $\mathcal{C}_{d}$ (see also [Fan_Lai:New_rational, Lemma 2.3]).
Remark 2.9.
In case we have an isometry of lattices $T\cong T^{\prime}$ and the signature of $T$ is of the form $(l,m)$ with $l\not=m$ then $T\not\cong T^{\prime}(-1)$. In particular this happens if $\operatorname{rk}T\not=4$ and the signature is $(2,\operatorname{rk}T-2)$.
3 Cubic fourfolds
Let $Y\subset\mathbb{P}^{5}_{\mathbb{C}}$ be a smooth cubic fourfold. In this case we have a polarized Hodge structure of K3 type $H^{4}(Y,\mathbb{Z})$ of weight $4$, which is a unimodular odd Hodge lattice of signature $(21,2)$. We denote by $\eta_{Y}\in A(Y)=H^{4}(Y,\mathbb{Z})\cap H^{2,2}(Y)$ the square of a hyperplane class, and let $H^{4}_{prim}(Y,\mathbb{Z})=\eta_{Y}^{\perp}$ be the primitive lattice of $Y$. As a lattice, $H^{4}_{prim}(Y,\mathbb{Z})$ is an even lattice isometric to $U^{\oplus 2}\oplus E_{8}^{\oplus 2}\oplus A_{2}$. For a very general cubic fourfold we have $A(Y)=\langle\eta_{Y}\rangle$. A cubic fourfold $Y$ is special if $\operatorname{rk}A(Y)\geq 2$, in this case it belongs to some Hassett divisor $\mathcal{C}_{d}\subset\mathcal{C}$ in the moduli space of cubic fourfolds with $d\equiv 0,2\mod 6$ and $d>6$, see [hassett2000special].
Lemma 3.1.
Let $Y$ and $Y^{\prime}$ be L-equivalent cubic fourfolds, then $\operatorname{disc}T(Y)=\operatorname{disc}T(Y^{\prime})$.
Proof.
This is 2.6 with the fact that $H^{4}(Y,\mathbb{Z})$ is unimodular for a cubic fourfold $Y$. ∎
Lemma 3.2.
Let $Y$ and $Y^{\prime}$ be L-equivalent cubic fourfolds such that $\operatorname{End}(T(Y))=\mathbb{Z}$ and $\operatorname{rk}T(Y)\not=4$, then there is a Hodge isometry $T(Y)\cong T(Y^{\prime})$.
Proof.
Since $Y$ and $Y^{\prime}$ are L-equivalent we have that $H^{4}(Y,\mathbb{Z})]=[H^{4}(Y^{\prime},\mathbb{Z})]\in\operatorname{HS}_{\mathbb{Z},4}$ and hence $T(Y)]=[T(Y^{\prime})]$. By 2.7 and 3.1 we know that there is a Hodge isometry $T(Y)\cong T(Y^{\prime})(\pm 1)$. The signature of $T(Y)$ is given by $(2,\operatorname{rk}T(Y)-2)$, and by 2.9 we have a Hodge isometry $T(Y)\cong T(Y^{\prime})$ since $\operatorname{rk}T(Y)\not=4$ . ∎
Theorem 3.3.
Let $Y\subset\mathbb{P}^{5}$ be a very general cubic fourfold. If $Y^{\prime}$ is a cubic fourfold which is L-equivalent to $Y$, then $Y\cong Y^{\prime}$. In particular, $Y$ and $Y^{\prime}$ are D-equivalent.
Proof.
Since $Y$ is very general, we have $T(Y)=H^{4}_{prim}(Y,\mathbb{Z})$ and in particular $\operatorname{rk}T(Y)=22$. Moreover, by 2.8 we have $\operatorname{End}(T(Y))=\mathbb{Z}$. In this case 3.2 gives a Hodge isometry $T(Y)\cong T(Y^{\prime})$. In particular, we have a Hodge isometry $H^{4}_{prim}(Y,\mathbb{Z})\cong H^{4}_{prim}(Y^{\prime},\mathbb{Z})$ and, up to composing it with $\mathrm{id}$, by 2.4 this extends to a Hodge isometry $H^{4}(Y,\mathbb{Z})\cong H^{4}(Y^{\prime},\mathbb{Z})$ mapping $\eta_{X}$ to $\eta_{Y^{\prime}}$. Hence, by the Torelli Theorem for cubic fourfolds [voisin1986theoreme] we can conclude that $Y\cong Y^{\prime}$. ∎
One might wonder if a similar statement holds for special subic fourfolds. This happens not to be the case, for this we recall a construction given by Fan-Lai [Fan_Lai:New_rational].
A very general cubic fourfold $Y\in\mathcal{C}_{20}$ contains a Veronese surface $V\subset\mathbb{P}^{5}$ and the system of quadrics containing $V$ determines a Cremona transformation
$F_{V}\colon\mathbb{P}^{5}\dashrightarrow\mathbb{P}^{5}.$
The indeterminacy locus of the birational inverse $F_{V}^{-1}$ is a Veronese surface $V^{\prime}\subset\mathbb{P}^{5}$ projective equivalent to $V$.
Proposition 3.4 ([Fan_Lai:New_rational, Theorem 1.1]).
Let $Y\in\mathcal{C}_{20}$ be a very general cubic fourfold containing a Veronese surface $V$ and let $Y^{\prime}$ be the proper image of $Y$ via $F_{V}$. Then $Y^{\prime}\in\mathcal{C}_{20}$ is a cubic fourfold containing $V^{\prime}$, and it is the unique cubic fourfold such that $Y$ and $Y^{\prime}$ are FM partners but $Y\not\cong Y^{\prime}$.
We recall that two cubic fourfolds $Y$ and $Y^{\prime}$ are called Fourier-Mukai (FM) partners if there is Fourier-Mukai equivalence $\mathcal{A}_{Y}\cong\mathcal{A}_{Y^{\prime}}$. Here $\mathcal{A}_{Y}:=\langle\mathcal{O},\mathcal{O}(1),\mathcal{O}(2)\rangle^{\perp}\subset D^{b}(Y)$, see [kuznetsov2009derived].
The previous examples are then easily seen to be L-equivalent.
Proposition 3.5.
There exist special cubic fourfolds $Y$ and $Y^{\prime}$ which are L-equivalent, but not isomorphic (and in particular not D-equivalent). Moreover, $Y$ is a FM partner of $Y^{\prime}$ and we can choose $Y$ so that $\operatorname{End}(T(Y))=\mathbb{Z}$.
Proof.
Let $Y\in\mathcal{C}_{20}$ be a very general cubic fourfold, containing a Veronese surface $V,$ and $Y^{\prime}$ be the proper image of the Cremona transformation $F_{V}$. The indeterminacy locus of $F_{V}$ is a Veronese surface $V\subset Y$ and the indeterminacy locus of the inverse $F_{V}^{-1}$ is a projective equivalent Veronese surface $V^{\prime}$. Thus, we can write the following relation
$Y]-[Y^{\prime}]=([Y\setminus V]+[V])-([Y^{\prime}\setminus V^{\prime}]+[V^{\prime}])=0\in K_{0}(\operatorname{Var}_{\mathbb{C}})$
since $V]=[V^{\prime}]$ and $Y\setminus V]=[Y^{\prime}\setminus V^{\prime}]$ from the above. In particular, $Y$ and $Y^{\prime}$ are L-equivalent and by 3.4 they are FM partners and not isomorphic. Since $Y$ is very general in $\mathcal{C}_{20}$, we have $\operatorname{End}(Y)=\mathbb{Z}$ by 2.8. ∎
For a cubic fourfold $Y$, we consider the lattice $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z}):=K_{top}(\mathcal{A}_{Y})$ which is isometric to $U^{\oplus 4}\oplus E_{8}(-1)^{\oplus 2}$. We denote by $A(\mathcal{A}_{Y})$ and $T(\mathcal{A}_{Y})$ the algebraic and transcendental sublattices, respectively. Moreover, $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ is endowed with a Hodge structure such that there is a Hodge isometry $T(Y)(-1)\cong T(\mathcal{A}_{Y})$, see [Addington_rationality].
Theorem 3.6.
Consider a very general cubic fourfold $Y\in\mathcal{C}_{d}$, for $d$ not divisible by $9$. If $Y^{\prime}$ is a cubic fourfold which is L-equivalent to $Y$, then $Y$ and $Y^{\prime}$ are FM partners.
Proof.
By [huybrechts2017k3, Theorem 1.5 iii)], this is equivalent to prove that there is a Hodge isometry $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})\cong\widetilde{H}(\mathcal{A}_{Y^{\prime}},\mathbb{Z})$.
Firstly, we observe that by [AT, Proposition 2.8] we have a Hodge isometry $T(Y)(-1)\cong T(\mathcal{A}_{Y})$. From [Fan_Lai:New_rational, Lemma 2.4] we know that the isometry class of $T(Y)$ is determined by $d$, and the discriminant group D(T(Y))D({T(Y)}) is cyclic when $d$ is not divided by $9$. Hence, in our hypothesis $D({T(\mathcal{A}_{Y})})$ is cyclic and also $D({A(\mathcal{A}_{Y})})$ is so from the fact that $D({T(\mathcal{A}_{Y})})$ is anti-isometric to $D({A(\mathcal{A}_{Y})})$ because $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ is unimodular.
Since $Y$ is very general in $\mathcal{C}_{d}$, we have that $\operatorname{End}(T(Y))=\mathbb{Z}$ by 2.8 and also $\operatorname{rk}T(Y)=21$. In this case 3.2 gives a Hodge isometry $T(Y)\cong T(Y^{\prime})$, in particular this implies that $D({A(\mathcal{A}_{Y})})$ and $D({A(\mathcal{A}_{Y^{\prime}})})$ are isometric. From the fact that $l(D({A(\mathcal{A}_{Y})}))=1$ and $A(\mathcal{A}_{Y})$ is indefinite with $\operatorname{rk}A(\mathcal{A}_{Y})=3$, we can apply 2.5 to get a Hodge isometry $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})\cong\widetilde{H}(\mathcal{A}_{Y^{\prime}},\mathbb{Z})$. ∎
Remark 3.7.
Note that by a conjecture of Huybrechts [huybrechts2023geometry, Conjecture 3.21] for two cubic fourfolds being FM partners implies being birational. A well known counter-example to the converse are two general Pfaffian cubic fourfolds, i.e. two general cubic fourfolds in $\mathcal{C}_{14}$, which are birational since both rational by [beauville1985variete] and in general not FM partners by [huybrechts2017k3, Theorem 1.5 iii)]. The same example shows that birationality does not imply L-equivalence for cubic fourfolds by virtue of Theorem 3.6.
4 Hyper-Kähler manifolds
A hyper-Kähler manifold $X$ will be a compact Kähler manifold which is simply connected and such that $H^{0}(X,\Omega^{2}_{X})=\sigma_{X}\cdot\mathbb{C}$ for an everywhere nondegenerate holomorphic 2-form $\sigma_{X}$. Since L-equivalence is for algebraic varieties, we will consider projective hyper-Kähler manifolds. In this case $H^{2}(X,\mathbb{Z})$ is a torsion-free polarized Hodge structure of K3 type of weight $2$, where the lattice structure is given by the Beauville-Bogomolov-Fujiki form of signature $(3,b_{2}(X)-3)$. The algebraic part coincides with the Néron-Severi lattice $\operatorname{NS}(X)=H^{1,1}(X)\cap H^{2}(X,\mathbb{Z})$ and it transcendental part is given by $T(X)=\operatorname{NS}(X)^{\perp}$, in analogy to the case of K3 surfaces (see for example [huybrechts2016lectures, Lemma 3.1]). Also, since $X$ is projective, the signature of $\operatorname{NS}(X)$ is of type (1,)(1,-) [gross2012calabi, Proposition 26.13]. One can refer for example to [gross2012calabi, debarre2022hyper] for generalities about hyper-Kähler manifolds.
Manifolds of K3[n] type are hyper-Kähler manifolds that are deformation equivalent to the Hilbert scheme of $n$ points on a K3 surface [beauville1983varietes]. Manifolds of OG10 type are hyper-Kähler manifolds that are deformation equivalent to the symplectic resolution of a moduli spaces of sheaves on a K3 surfaces as in [OG10].
Lemma 4.1.
Let $X$ and $X^{\prime}$ be L-equivalent projective hyper-Kähler manifolds of the same deformation type such that $\operatorname{disc}H^{2}(X,\mathbb{Z})=p$ is a prime, then $\operatorname{disc}T(X)=c\cdot\operatorname{disc}T(X^{\prime})$ with $c\in\{1,p,\frac{1}{p}\}$.
Proof.
We have $m=|G(X)|=|G(X^{\prime})|$, moreover $m|\operatorname{disc}T(X)$ and $m|\operatorname{disc}\operatorname{NS}(X)$ and similarly for $X^{\prime}$. Observing that from (1)
$pm^{2}=\operatorname{disc}T(X)\cdot\operatorname{disc}\operatorname{NS}(X)$
we deduce that $\operatorname{disc}T(X)\in\{m,pm\}$, similarly for $X^{\prime}$. ∎
Lemma 4.2.
Let $X$ and $X^{\prime}$ be L-equivalent projective hyper-Kähler manifolds of the same deformation type and assume that $\operatorname{disc}H^{2}(X,\mathbb{Z})=p$ a prime number, $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$. If there is a primitive embedding $H^{2}(X,\mathbb{Z})\subseteq\Lambda$ in a unimodular lattice $\Lambda$ with $2\operatorname{rk}T(X)>\operatorname{rk}\Lambda$, then there is a Hodge isometry $T(X)\cong T(X^{\prime})$.
Proof.
Since $\operatorname{End}(T(M))=\mathbb{Z}$, then by 2.7 together with 4.1 there is a Hodge isometry $T(X)\cong T(X^{\prime})(q)$ for $q\in\{\pm 1,\pm p,\pm\frac{1}{p}\}$. Up to exchanging $X$ and $X^{\prime}$, we can assume $q\in\mathbb{Z}$. From the hypothesis that there is a primitive embedding $H^{2}(X,\mathbb{Z})\subseteq\Lambda$ in a unimodular lattice $\Lambda$ and $2\operatorname{rk}T(X)>\operatorname{rk}\Lambda$, we can apply 2.1 to deduce that $q=\pm 1$. Since $X$ and $X^{\prime}$ are projective, then $T(X)$ and $T(X^{\prime})$ have signature (2,)(2,-) and by 2.9 with $\operatorname{rk}T(X)\not=4$ we get a Hodge isometry $T(X)\cong T(X^{\prime})$. ∎
Remark 4.3.
We recall the following:
  1. 1.
    For $X$ of OG10 type, we have that $H^{2}(X,\mathbb{Z})\cong U^{\oplus 3}\oplus E_{8}(-1)^{\oplus 2}\oplus A_{2}(-1)$ as lattices. In particular, $\operatorname{disc}H^{2}(X,\mathbb{Z})=3$ and there is a primitive embedding in the unimodular lattice $\Lambda=U^{\oplus 5}\oplus E_{8}(-1)^{\oplus 2}$ of rank $26$. Hence, 4.2 can be applied for $\operatorname{rk}T(X)>13$.
  2. 2.
    For $X$ of K3[2] type, we have that $H^{2}(X,\mathbb{Z})\cong U^{\oplus 3}\oplus E_{8}(-1)^{\oplus 2}\oplus[-2]$ as lattices. In particular, $\operatorname{disc}H^{2}(X,\mathbb{Z})=2$ and there is a primitive embedding in the unimodular lattice $\Lambda=U^{\oplus 4}\oplus E_{8}(-1)^{\oplus 2}$ of rank $24$. Hence, 4.2 can be applied for $\operatorname{rk}T(X)>12$.
4.1 Hyper-Kähler manifolds from a cubic fourfold
4.1.1 The Fano variety of lines
We firstly consider the Fano variety of lines $F(Y)$ associated to a cubic fourfold $Y$, this is a projective hyper-Kähler manifold of K3[2] type [beauville1985variete].
Lemma 4.4.
Let $Y$ and $Y^{\prime}$ be L-equivalent cubic fourfolds, then their Fano varieties of lines $F(Y)$ and $F(Y^{\prime})$ are L-equivalent.
Proof.
We know by [okawa2021example, Lemma 2.1] that $Y^{[2]}$ and $Y^{\prime[2]}$ are L-equivalent. Moreover, by [galkin2014fano, Theorem 5.1] we have that
$\mathbb{L}^{2}([F(Y)]-[F(Y^{\prime})])=([Y^{[2]}]-[Y^{\prime[2]}])+[\mathbb{P}^{4}]([Y]-[Y^{\prime}])\in K_{0}(\operatorname{Var}_{\mathbb{C}}).$
The first part of the statement follows multiplying by $\mathbb{L}^{m}$ where $m$ is big enough so that $Y^{[2]}]-[Y^{\prime[2]}]$ and $Y]-[Y^{\prime}]$ are annihilated, recalling that the Grothendiek ring of varieties is commutative. ∎
The following gives more examples of projective hyper-Kähler manifolds that are L-equivalent, but not birational. The first such examples were constructed in [okawa2021example], here we present some cases that were not included and that do not have an associate K3 surface (see [hassett2000special]). The previous examples are also known to be D-equivalent, we expect our examples to be D-equivalent as well in view of [meinsma2025counterexamples, Conjecture 1.5].
Proposition 4.5.
Let $Y\in\mathcal{C}_{20}$ be a very general cubic fourfold, and let $Y^{\prime}\in\mathcal{C}_{20}$ be as in 3.4. Then the Fano varieties of lines $F(Y)$ and $F(Y^{\prime})$ are L-equivalent, but not birational.
Proof.
Firstly, we know that $F(Y)$ and $F(Y^{\prime})$ are L-equivalent by 4.4. We also recall that there is a Hodge isometry $H^{4}_{prim}(Y,\mathbb{Z})\cong H^{2}_{prim}(F(Y),\mathbb{Z})(-1)$, where $H^{2}_{prim}(F(Y),\mathbb{Z})$ is the orthogonal of the polarization $H$ of $F(Y)$, and $H^{2}=6$, from [beauville1985variete].
Assume $F(Y)$ is birational to $F(Y^{\prime})$. Then by the Torelli Theorem for hyper-Kähler manifold [markman2011survey, Theorem 1.3] there exists a Hodge isometry $H^{2}(F(Y),\mathbb{Z})\cong H^{2}(F(Y^{\prime}),\mathbb{Z})$, which might a priori not preserve the polarization. The algebraic lattice of a general cubic $Y\in\mathcal{C}_{20}$ is given by
$A(Y)=\langle\eta_{Y},V\rangle\cong\begin{bmatrix}3&4\\ 4&12\end{bmatrix}$
where $V\subset Y$ is a Veronese surface, see [Fan_Lai:New_rational] for example. The algebraic primitive lattice is then given by $A_{prim}(Y)=\eta_{Y}^{\perp A(Y)}=\langle-4\eta_{X}+3V\rangle\cong[60]$. Let $H$ be the polarization of $F(Y)$, then $\operatorname{NS}(F(Y))$ is an overlattice of $\langle H\rangle\oplus\langle\alpha\rangle$ with $H^{2}=6$ and $\alpha^{2}=-60$. Since $\operatorname{disc}T(Y)=\operatorname{disc}T(F(Y))=20$, the index of the inclusion $\langle H\rangle\oplus\langle\alpha\rangle\subset\operatorname{NS}(F(Y))$ must be $3$. This shows that
$\operatorname{NS}(F(Y))=\langle H,\frac{H-\alpha}{3}\rangle\cong\begin{bmatrix}6&2\\ 2&-6\end{bmatrix}$
and one can check that $O(\operatorname{NS}(F(Y))\cong\{\pm\mathrm{id}\}$. In particular, up to composing with $\mathrm{id}$, twe find a Hodge isometry $H^{2}(F(Y),\mathbb{Z})\cong H^{2}(F(Y^{\prime}),\mathbb{Z})$ preserving the polarization and this induces a Hodge isometry $H_{prim}^{4}(Y,\mathbb{Z})\cong H_{prim}^{4}(Y^{\prime},\mathbb{Z})$. However, this is not possible since $Y\not\cong Y^{\prime}$ and we conclude that $F(Y)$ and $F(Y^{\prime})$ are not birational. ∎
4.1.2 The LSV manifolds
We now treat the case of projective hyper-Kähler manifolds that are birational to the Laza–Saccà–Voisin (LSV) manifold $J(Y)$ (or its twisted version $J^{t}(Y)$) for a cubic fourfold $Y$. These are hyper-Kähler manifolds of OG10 type obtained as compactifications of the fibration (twisted) intermediate Jacobian of hyperplane section of cubic fourfolds, we refer to [laza2017hyper, voisin2016hyper] for their construction.
Proposition 4.6.
Let $X$ be a manifold of OG10 type birational to $J(Y)$ (or $J^{t}(Y)$) for a very general cubic fourfold $Y$. Let $T^{\prime}$ be a cubic fourfold and $X^{\prime}$ be another manifold of OG10 type birational to $J(Y^{\prime})$ (or $J^{t}(Y^{\prime})$). If $X$ is L-equivalent to $X$, then $Y\cong Y^{\prime}$ and $X$ is birational to $X^{\prime}$.
Proof.
We have that $T(X)$ is Hodge anti-isometric to $T(Y)$, see [laza2017hyper, sacca2023birational] and [billi2024non, Proposition 5.1]. Since $Y$ is very general we know by 2.8 that $\operatorname{End}(T(X))=\mathbb{Z}$. In this case we have $\operatorname{rk}T(X)=22$ and by 4.3 we can apply 4.2 to deduce that there is a Hodge isometry $T(X)\cong T(X^{\prime})$. In particular, we obtain a Hodge isometry $T(Y)\cong T(Y^{\prime})$ and by the Torelli Theorem for cubic fourfold [voisin1986theoreme] with the fact that $Y$ and $Y^{\prime}$ are very general, we get $Y\cong Y^{\prime}$. In conclusion, we have that $X$ and $X^{\prime}$ are birational to $J(Y)$ (or $J^{t}(Y)$). ∎
We have that if $Y$ is very general then $\operatorname{NS}(J(Y))\cong U$ and $\operatorname{NS}(J^{t}(Y))\cong U(3)$, moreover there are rational Hodge isometries $T_{\mathbb{Q}}(J(Y))\cong T_{\mathbb{Q}}(J^{t}(Y))\cong T_{\mathbb{Q}}(Y)(-1)$. This implies that there exists $q\in\mathbb{Q}$ such that there is a Hodge isometry $T(J(Y))\cong T(J^{t}(Y))(q)$, since there is an isometry $T(J(Y))\cong T(J^{t}(Y))\cong U^{\oplus 2}\oplus E_{8}(-1)^{\oplus 2}\oplus A_{2}(-1)$ we then have a Hodge isometry $T(J(Y))\cong T(J^{t}(Y))$.
We then wonder if $J(Y)$ and $J^{t}(Y)$ can be L-equivalent, and notice that by [Sacca_birationality_twisted, Corollary 3.10] they are not birational.
4.1.3 Moduli spaces on the Kuznetsov component
We finally treat the case of hyper-Kähler manifolds that are constructed from moduli spaces of Bridgeland (semi-)stable objects in the Kuznetsov component $\mathcal{A}_{Y}$ of a cubic fourfold $Y$.
The first stability condition on $\mathcal{A}_{Y}$ was constructed in [BLMS]. We distinguish the connected component $\operatorname{Stab}^{\dagger}(\mathcal{A}_{Y})\subset\operatorname{Stab}(\mathcal{A}_{Y})$ of the manifold of stability conditions containing such a condition, and for a Mukai vector $v\in\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ and a stability condition $\sigma\in\operatorname{Stab}^{\dagger}(\mathcal{A}_{Y})$, denote by $M_{v}(\sigma)$ the moduli space of $\sigma$-semistable objects in $\mathcal{A}_{Y}$ with Mukai vector $v\in A(\mathcal{A}_{Y})$. We also recall that there are always two algebraic classes $\lambda_{1},\lambda_{2}\in A(\mathcal{A}_{Y})$ spanning a primitive sublattice isometric to $A_{2}$, and when $Y$ is very general we have $A(\mathcal{A}_{Y})\cong A_{2}$.
Proposition 4.7 ([BLMNPS, LPZ_OG10]).
Let $v\in\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ and $\sigma\in\operatorname{Stab}^{\dagger}(\mathcal{A}_{Y})$ be a $v$-generic stability condition.
  1. 1.
    If $v$ is primitive, then $M_{v}(\sigma)$ is hyper-Kähler of K3[n] type with $n=\frac{v^{2}+2}{2}$.
  2. 2.
    If $v=2v_{0}$ with $v_{0}$ primitive and $v_{0}^{2}=2$, then $M_{v}(\sigma)$ admits a symplectic resolution $\widetilde{M}_{v}(\sigma)$ via blow-up which is a hyperkähler manifold of OG10 type.
For this reason, we will call a vector $v\in\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ of O’Grady type if $v=2v_{0}$ for a primitive vector $v_{0}$ with $v_{0}^{2}=2$.
Most of the propositions proved in the previous sections rely on finding some connection between the L-equivalence and Hodge theoretic properties of cubic fourfolds. This is why the following lemma is particularly useful in order to deal with $M_{v}(\sigma)$ for some $\sigma$ in $\operatorname{Stab}^{\dagger}(\mathcal{A}_{Y})$ and $v\in\widetilde{H}(Y,\mathbb{Z})$.
Lemma 4.8.
Let $\sigma$ be a $v$-generic stability condition in $\operatorname{Stab}^{\dagger}(\mathcal{A}_{Y})$ for either a primitive or a O’Grady type Mukai vector $v\in\widetilde{H}(Y,\mathbb{Z})$. Then $T(M_{v}(\sigma))\simeq T(Y)(-1)$.
Proof.
We know by [GGO, Proposition 2.8] and [BLMNPS, Theorem 1.6] that in both cases there exists a Hodge isometry $H^{2}(M_{v}(\sigma),\mathbb{Z})\simeq v^{\perp}\subset\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$. Moreover, by [AT, Proposition 2.8] there exists another Hodge isometry $\langle\lambda_{1},\lambda_{2}\rangle^{\perp}=A_{2}^{\perp}\simeq H^{4}_{prim}(Y,\mathbb{Z})(-1)$ where the first orthogonal complement is taken in the Mukai lattice of $\mathcal{A}_{Y}$. As $v\in\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ lies in the algebraic part we deduce from both the isometries that $T(M_{v}(\sigma))\simeq T(Y)(-1)$. In case $v$ is of O’Grady type, then $\widetilde{M}_{v}(\sigma)$ is given by a blow-up on the singular locus of $M_{v}(\sigma)$, so that $T(\widetilde{M}_{v}(\sigma))\simeq T(M_{v}(\sigma))\simeq T(Y)(-1)$. ∎
The following lemma tells us that when $Y$ is very general, a primitive Mukai vector $v\in A(\mathcal{A}_{Y})$ is determined by its square up to isometry.
Lemma 4.9.
Let $v,w\in A_{2}$ be two primitive elements with the same square. Then $v^{\perp}\simeq w^{\perp}$ as lattices.
Proof.
Let $v=ae_{1}+be_{2}$ for a basis $\{e_{1},e_{2}\}$ of $A_{2}$. Then a generator of $v^{\perp}$ will be a primitive element $u=xe_{1}+ye_{2}$ such that (2ab)x+(2ba)y=0(2a-b)x+(2b-a)y=0. We call p:=(2ab)p:=(2a-b) and q:=(2ba)q:=(2b-a). The diophantine equation $px+qy=0$ has solutions
$x=k\frac{q}{g},\ y=-k\frac{p}{g},\qquad g:=\gcd(p,q),\quad k\in\mathbb{Z}.$
The only primitive solutions are thus given by $k=\pm 1$. In both cases we get
$u^{2}=\frac{2}{g^{2}}\left(q^{2}+p^{2}+pq\right)=\frac{6}{g^{2}}\left(a^{2}+b^{2}-ab\right)=\frac{3}{g^{2}}v^{2}.$
Note that $p\equiv q\equiv 2(a+b)\mod 3$. Moreover, $g\mid(2p+q)=3a$ and $g\mid(2q+p)=3b$. As $v$ is primitive we know that $\gcd(a,b)=1$, therefore $g\mid 3$, i.e. $g\in\left\{\pm 1,\pm 3\right\}$. Now, we have the identity $a^{2}+b^{2}-ab\equiv(a+b)^{2}\mod 3$. Therefore,
$0\equiv v^{2}=2(a^{2}+b^{2}-ab)\mod 3\iff 0\equiv a+b\equiv p\equiv q\mod 3\iff g=\pm 3.$
This shows that the square of the primitive element $u$ that generates the orthogonal complement of a primitive vector $v\in A_{2}$ is $u^{2}=\frac{v^{2}}{3}$ if $v^{2}\equiv 0\mod 3$ or, else, $u^{2}=3v^{2}$. In both cases it is uniquely determined by the square of $v$. ∎
Remark 4.10.
If $v=kv_{0}$ for $k\in\mathbb{Z}$ and $v_{0}$ primitive we get that an element $u\in v^{\perp}\iff u\in v_{0}^{\perp}$, so the same result of 4.9 holds for $v=kv_{0}$, $w=kw_{0}$ with $v_{0},w_{0}$ primitive vectors with the same square.
Now we are ready to state and prove the following proposition in the same fashion of 4.6.
Proposition 4.11.
Let $Y$ and $Y^{\prime}$ be two cubic fourfolds and assume that $Y$ is very general. Let $v\in A(\mathcal{A}_{Y})$ and $w\in A(\mathcal{A}_{Y^{\prime}})$ be either primitive with same square or of O’Grady type. Let $\sigma_{Y}\in\operatorname{Stab}^{\dagger}(Y)$ and $\sigma_{Y^{\prime}}\in\operatorname{Stab}^{\dagger}(Y^{\prime})$ be a $v$-generic and a $w$-generic stability condition, respectively. Suppose that $M_{v}(\sigma_{Y})$ and $M_{w}(\sigma_{Y^{\prime}})$ are L-equivalent, then $Y\cong Y^{\prime}$ and $M_{v}(\sigma_{Y})$ is birational to $M_{w}(\sigma_{Y^{\prime}})$.
Proof.
By 4.8 we get that $T(M_{v}(\sigma_{Y}))$ is Hodge anti-isometric to $T(Y)$ and $T(M_{w}(\sigma_{Y}))$ to $T(Y^{\prime})$. Since $Y$ is very general we know by 2.8 that $\operatorname{End}(T(X))=\mathbb{Z}$. Since $Y$ is very general, we have $T(Y)=H_{prim}^{4}(Y,\mathbb{Z})$ and similarly for $Y^{\prime}$. From the fact that $T(Y)$ and $T(Y^{\prime})$ are isometric of signature $(22,2)$, by 2.7 and 2.9 we get Hodge isometries $T(Y)\simeq T(M_{v}(\sigma_{Y}))(-1)\simeq T(M_{w}(\sigma_{Y^{\prime}}))(-1)\simeq T(Y^{\prime})$. From the first and the last Hodge isometries we get a Hodge isometry $H^{4}(Y,\mathbb{Z})\cong H^{4}(Y^{\prime},\mathbb{Z})$. The Torelli theorem for cubic fourfolds implies then then $Y\simeq Y^{\prime}$. Moreover, we have $A(\mathcal{A}_{Y})\simeq A(\mathcal{A}_{Y^{\prime}})\simeq A_{2}$ and by [AT, Proposition 2.8] we get Hodge isometries $A_{2}^{\perp}=T(\mathcal{A}_{Y})\cong H^{4}_{prim}(Y,\mathbb{Z})(-1)\simeq H^{4}_{prim}(Y^{\prime},\mathbb{Z})(-1)\cong T(\mathcal{A}_{Y^{\prime}})$.
As $v,w\in A_{2}$ share the same square and are either primitive or of O’Grady type we deduce from 4.9, with the fact that $O(A_{2})\to O(D({A_{2}}))$ is surjective, that they have isomorphic embeddings in $A_{2}$. So, we can deduce the following chain of Hodge isometries
$H^{2}(M_{v}(\sigma_{Y}),\mathbb{Z})\simeq v^{\perp}\simeq w^{\perp}\simeq H^{2}(M_{w}(\sigma_{Y^{\prime}}),\mathbb{Z}).$ (2)
If $v$ and $w$ are of O’Grady type then the Hodge isometry of (2) induces a Hodge isometry $H^{2}(\widetilde{M}_{v}(\sigma_{Y}))\simeq H^{2}(\widetilde{M}_{w}(\sigma_{Y^{\prime}}))$. Moreover, by [onorati_monodromy] the monodromy group of a manifolds of OG10 type $X$ is given by $O^{+}(H^{2}(X,\mathbb{Z}))$, thus, up to composing by $\mathrm{id}$ we can assume that the Hodge isometry is monodromy. The Hodge theoretic Torelli theorem for hyper-Kähler manifolds [markman2011survey, Theorem 1.3] implies that $M_{v}(\sigma_{Y})$ is birational to $M_{w}(\sigma_{Y^{\prime}})$.
If $v$ and $w$ are primitive, then $M_{v}(\sigma_{Y})$ and $M_{w}(\sigma_{Y^{\prime}})$ are of type $K3^{[n]}$ with $n=\frac{v^{2}+2}{2}$. By the previous discussion, the Hodge isometry $T(\mathcal{A}_{Y})\cong T(\mathcal{A}_{Y^{\prime}})$ extends to a Hodge isometry of the Mukai lattices $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})\cong\widetilde{H}(\mathcal{A}_{Y^{\prime}},\mathbb{Z})$ restricting to (2) with the opportune identifications, and hence [markman2011survey, Corollary 9.9] yields that $M_{v}(\sigma_{Y})$ is birational to $M_{w}(\sigma_{Y^{\prime}})$. ∎
Remark 4.12.
By [LPZ_OG10, Proposition 6.7] the OG10 manifold $\widetilde{M}_{v}(\sigma)$ for $v$ of O’Grady type provides a compactification of the Lagrangian fibration of twisted intermediate Jacobians. Thus, $\widetilde{M}_{v}(\sigma)$ is birational to the twisted LSV manifold $J^{t}(Y)$ [voisin2016hyper]. In particular, when $Y$ is very general, for this choice of $v$, we have that 4.11 and 4.6 are equivalent.
The tight relation between the Hodge structures of $M_{v}(\sigma)$ and the underlying cubic fourfold $Y$ yields also the following.
Corollary 4.13.
Suppose that in the hypothesis of 4.11 we ask $Y$ to be very general in $\mathcal{C}_{d}$ for $d$ not divisible by 9. Then $Y$ and $Y^{\prime}$ are FM partners.
Proof.
The proof of the fact that $Y$ and $Y^{\prime}$ are FM partners is the same of Theorem 3.6, once we note that by 2.7 we have that $T(Y)(-1)\simeq T(M_{v}(\sigma))\simeq T(\mathcal{A}_{Y})$. ∎
However, in general we do not get birationality of the moduli spaces in this case as shown in 4.5.
4.2 Hyper-Kähler manifolds from K3 surfaces
4.2.1 Desingularizations of moduli spaces of sheaves on a K3
We firstly consider the case where the hyper-Kähler manifolds are birational to desingularizations of moduli spaces of sheaves on K3 surfaces.
Proposition 4.14.
Let $X$ be a manifold of OG10 type with $\operatorname{NS}(X)=U\oplus A_{2}(-1)$, $\operatorname{disc}T(X)=1$ and $\operatorname{End}(T(X))=\mathbb{Z}$. Let $X^{\prime}$ be another manifold of OG10 type which is L-equivalent to $X$, then $X$ and $X^{\prime}$ are birational. Moreover, $X$ and $X^{\prime}$ are D-equivalent.
Proof.
Since $\operatorname{disc}T(X)=1$ there is a primitive vector $\sigma\in\operatorname{NS}(X)$ such that $\sigma^{2}=-6$ of divisibility $3$ in $H^{2}(X,\mathbb{Z})$. By [FGG:OG10_as_moduli, Theorem 3.13] we have that $X$ is birational to a moduli space of sheaves on a K3 surface. Moreover, we have $\operatorname{rk}T(X)=20>13$ and by 4.2 there is a Hodge isometry $T(X)\cong T(X^{\prime})$ in view of 4.3. In particular, $T(X^{\prime})$ is also unimodular and from this it follows that $\operatorname{NS}(X^{\prime})$ has the same discriminant form as $\operatorname{NS}(X)$ (which is the one of $H^{2}(X,\mathbb{Z})$), hence $\operatorname{NS}(X)$ and $\operatorname{NS}(X^{\prime})$ are in the same genus. Since the two lattices are even indefinite and $\operatorname{rk}\operatorname{NS}(X)\geq l(\operatorname{NS}(X))+2$, we can extend the Hodge isometry $T(X)\cong T(X^{\prime})$ to a Hodge isometry $H^{2}(X,\mathbb{Z})\cong H^{2}(X^{\prime},\mathbb{Z})$ by 2.5. Up to composing it with $\mathrm{id}$, the Hodge isometry can be taken to be monodromy by [onorati_monodromy, Theorem 5.4] and the Torelli Theorem for hyper-Kähler manifolds [markman2011survey, Theorem 1.3] implies that $X$ and $X^{\prime}$ are birational.
In particular, they are birational to a moduli space of sheaves on a K3 surface and then by [halpern2020derived, Theorem 4.5.1] they are D-equivalent.
4.2.2 Hilbert schemes of points on a K3 surface
Let $X$ be hyper-Kähler of K3[n] type, in [Taelman:derived_equivalences, beckmann2023derived] the authors define a Hodge lattice $\Lambda_{X}\subset\widetilde{H}(X,\mathbb{Q})$. As abstract lattices, $\Lambda_{X}\cong U\oplus H^{2}(X,\mathbb{Z})$ and its Hodge structure is such that $H^{2}(X,\mathbb{Z})^{\perp}\subseteq\Lambda_{X}^{1,1}$. If two manifolds of K3[n] type are D-equivalent, then there is a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$ by [beckmann2023derived, Theorem 9.2].
Lemma 4.15.
Let $X$ be a manifold of K3[n] type such that $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$. Assume that $X^{\prime}$ is another manifold of K3[n] type L-equivalent to $X$ with $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})$. Assume that $D({\operatorname{NS}(X)})$ and $D({\operatorname{NS}(X^{\prime})})$ are isometric, then there is a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$.
Proof.
Since $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})$, $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$, then by 2.7 and 2.9 we have a Hodge isometry $T(X)\cong T(X^{\prime})$. From the fact that $U=H^{2}(X,\mathbb{Z})^{\perp}\subseteq\Lambda_{X}^{1,1}$ we have $\operatorname{rk}\Lambda_{X}^{1,1}\geq l(\Lambda_{X}^{1,1})+2$. Moreover, since $\operatorname{NS}(X)$ and $\operatorname{NS}(X^{\prime})$ have the same genus then also $\Lambda_{X}^{1,1}$ and $\Lambda_{X^{\prime}}^{1,1}$ have the same genus. We are then in the place to apply 2.5 to get a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$.
Moreover, in some specific cases there is a better description of Hodge isometries induced by derived equivalence.
Proposition 4.16.
Let $X=S^{[n]}$ for n1n-1 a prime power and a K3 surface $S$ with $U\subseteq\operatorname{NS}(S)$, and assume that $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$. If $X^{\prime}$ is a manifold of K3[n] type which is L-equivalent to $X$ and $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})=m$ is coprime to 2(n1)2(n-1), then $X$ and $X^{\prime}$ are D-equivalent.
Proof.
In this case we know from [beckmann2023derived, Theorem 9.8] that any Hodge isometry $\Lambda_{X}\cong\Lambda_{X}$ preserving the orientation of a positive-definite real $4$-plane is induced by a derived equivalence $D^{b}(X)\cong D^{b}(X^{\prime})$. It suffices then to show that 4.15 can be applied in this situation giving a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$ and that this Hodge isometry can be chosen to be orientation-preserving.
Firstly, we need to show that $D({\operatorname{NS}(X)})$ is isometric to $D({\operatorname{NS}(X^{\prime})})$. We apply 2.7 to obtain a Hodge isometry $T(X)\cong T(X^{\prime})$. By hypothesis, the discriminant group $D({T(X^{\prime})})$ does not contain any subgroup isomorphic to $\mathbb{Z}/2(n-1)\mathbb{Z}$ so that the primitive embedding $T(X^{\prime})\subset H^{2}(X^{\prime},\mathbb{Z})$ has trivial embedding subgroup and then the discriminant of the orthogonal complement $\operatorname{NS}(X^{\prime})$ is given by $D({\operatorname{NS}(X^{\prime})})\cong D({T(X^{\prime})})\oplus D({[-2(n-1)]})$, which is isometric to the discriminant form of $\operatorname{NS}(X)$. We are now in the hypothesis of 4.15, which gives an Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$.
We finally observe that the Hodge isometry can be chosen to be orientation preserving. In fact, from $U\subseteq\operatorname{NS}(X)$ we have that $\operatorname{rk}\Lambda_{X}^{1,1}\geq l(D({\Lambda_{X}^{1,1}}))+4$ so that by [Nikulin_integral, Corollary 1.13.5] the lattice $\Lambda_{X}^{1,1}$ splits a copy of $U$. Hence, if the Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$ is reversing the orientation, after composing with the isometry acting as $\mathrm{id}$ on the copy of $U$ and trivially on the complement, we obtain a Hodge isometry preserving the orientation. ∎